OpenStudy (anonymous):

evaluate if x= -10 and y=-3 -4y^2

OpenStudy (anonymous):

144

OpenStudy (radar):

Where is the x?

OpenStudy (anonymous):

he has a series of exercises it is just one of them and in that one does not have a x on it

OpenStudy (anonymous):

yup

OpenStudy (aroub):

its -144 or 144

OpenStudy (anonymous):

-4 * -3 is 12

OpenStudy (radar):

OK so substitute that value\[-4(-3)^{2}\] -4(9)=-36

OpenStudy (aroub):

ok it 12 squared

OpenStudy (anonymous):

i like radar answer better than mine

OpenStudy (aroub):

yes i think so

OpenStudy (anonymous):

what about this one

OpenStudy (radar):

which one?

OpenStudy (anonymous):

-xy^3

OpenStudy (radar):

what do you want to evaluate? x=?? y=???

OpenStudy (aroub):

i think both

OpenStudy (aroub):

right?

OpenStudy (radar):

\[-xy ^{3}=-(-10)(-3)^{3}=-(-10)(-27)=-(270)=-270\]

OpenStudy (radar):

was my assumption correct? or are the evaluators somethin other than -10 and -3?

OpenStudy (anonymous):

there still the same

OpenStudy (radar):

O.K. did you follow the solution from step to step?

OpenStudy (anonymous):

yea i got it

OpenStudy (radar):

Excellent, don't rush, be neat on your paper, and good luck.

OpenStudy (anonymous):

x - y^3

OpenStudy (radar):

\[x-y ^{3}=-10-(-3)^{3}=-10-(-27)=-10+27=17\]

OpenStudy (anonymous):

your the best!!!!!!!

OpenStudy (radar):

Order of operations require doing the exponents before adding or multiplying/dividing.

OpenStudy (radar):

Good luck with it, take your time and do it step by step following the rules of order of operations. Also keep good track of signs (which is my weakness)